the Hamilton corporation has 6 millions shares of stock outstanding and will report earning of $6650000 in the current year. The company is considering the insurance of 1 million shares which can only be issued at $30 per share. A. what is the income per share? B. earning per share after addition Income?
Answer #1

HI

Here stock outstanding = 6 million

earning = $6650000

A) current income per share= Earning/ stock outstanding = 6.65 / 6 = $1.11

B) after issuing new stocks total stock outstanding = 6+1 =7 million

it will not impact net income

Hence new Earning per share = 6.65/7 = $0.95

Thanks